OpenStudy (anonymous):

√(x+2)+4=x Help, please! This is my last problem!!

OpenStudy (anonymous):

\[\LARGE \sqrt{x+2}+4=x\]

OpenStudy (anonymous):

start with \[\sqrt{x+2}=x-4\] square both sides to get \[x+2=(x-4)^2\]\[x+2=x^2-8x+16\]\[x^2-9x+14=0\] then factor

OpenStudy (anonymous):

\((x-7)(x-2)=0\) so either \(x=7\) or \(x=2\) now check to see if they work

OpenStudy (anonymous):

7 works because \[\sqrt{7+2}+4=7\] but 2 does not

OpenStudy (anonymous):

Thank you :D